(x^2-x-2)y’+3xy=x^2-4x+4
1
Expert's answer
2020-10-13T17:30:53-0400

The given D.E is a lin ear 1st order differential equation:

"(x^2-x-2)\frac{dy}{dx}+3xy =( x^2-4x+4)"



Now this can be rewritten as

"\frac{dy}{dx}+\frac{3x}{(x^2-x-2)}*y=\frac{(x^2-4x+4)}{(x^2-x-2)}"

"or, \frac{dy}{dx}+\frac{3x}{(x+1)(x-2)}*y=\frac{(x-2)^2}{(x+1)(x-2)}\\or, \frac{dy}{dx}+\frac{3x}{(x+1)(x-2)}*y=\frac{(x-2)}{(x+1)}"


Calculating INTEGRATING FACTOR (I.F)


"I.F= e^{\int {\frac{3x}{(x+1)(x-2)}}dx}"

"Now \int {\frac{3x}{(x+1)(x-2)}}dx = \int \{\frac{1}{x+1}+\frac{2}{x-2}\}dx"

"=\int \frac{1}{x+1}dx + 2\int \frac{1}{x+2}dx"

"= log|x+1|+2log|x+2|"

"= log\{(x+1)(x-2)^2\}"

"\therefore I.F = e^{log\{(x+1)(x-2)^2\}}=(x+1)(x-2)^2"


The solution to the given D.E is


"y*I.F=\int I.F*\frac{(x-2)}{(x+1)}dx + C"

"\therefore y* (x+1)(x-2)^2= \int \{(x+1)(x-2)^2\frac{(x-2)}{(x+1)}\}dx +C\\or, y* (x+1)(x-2)^2= \int (x-2)^3dx+C"

"or,y* (x+1)(x-2)^2= \frac{(x-2)^4}{4}+C"


The required solution is

"y* (x+1)(x-2)^2= \frac{(x-2)^4}{4}+C"


ANSWER